OpenStudy (moonlitfate):

The radius of a right circular cylinder is (3t +4)^(1/2) and its height is t^6 , where t is time in seconds and the dimensions are in inches. Find the rate of change of the volume of the cylinder, V,with respect to time, including the appropriate units.

OpenStudy (moonlitfate):

I know that the volume of a right cylinder is: V= pi*r^2*h

OpenStudy (anonymous):

You'll get V in terms of t, and then differentiate it with respect to t

OpenStudy (moonlitfate):

So plug in the give radius and height for r and h in the equation for volume, right?

OpenStudy (moonlitfate):

\[V = \pi(\sqrt{3t + 4})(t^{6})\]

OpenStudy (moonlitfate):

And I know that to differentiate it that you have to use the chain rule, but I've only done that with two functions. How would you do it in this case since there are 3?

OpenStudy (dan815):

its like this